OpenStudy (anonymous):

trig integral 1/sqrt. (sin2x) dx

OpenStudy (xapproachesinfinity):

hmmm \(\large \int\frac{1}{\sqrt{sin2x}}dx\) is it this?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I don't think there's an elementary result...

OpenStudy (xapproachesinfinity):

looks like a dirty integral!

OpenStudy (anonymous):

sorry my mistake..... its integral of sqrt. (1-cos4x) dx

OpenStudy (xapproachesinfinity):

\(1-cos4x ) right?

OpenStudy (xapproachesinfinity):

\(1-cos4x \)

OpenStudy (anonymous):

with a square root

OpenStudy (xapproachesinfinity):

yeah! im just making sure it is not a power! use double angle identity first

OpenStudy (xapproachesinfinity):

cos(2x)=2cos^2x-1 so cos4x=2cos^2(2x)-1

OpenStudy (xapproachesinfinity):

with this you should get rid of square root

OpenStudy (anonymous):

will it be sqrt 2 -sqrt (2 cos^2 (2x))?

OpenStudy (xapproachesinfinity):

hmmm \(1-cos4x=1-(2cos^22x-1)=2-2cos^22x=2(1-cos^22x)=2sin^22x\)

OpenStudy (xapproachesinfinity):

putting this back into our integral \(\large \int\frac{1}{\sqrt{2sin^22x}}dx\)

OpenStudy (xapproachesinfinity):

since we achieved this \(\large \int \frac{1}{\sqrt{2}sin(2x)}dx=\frac{1}{\sqrt{2}}\int\frac{1}{sin(2x)}dx\)

OpenStudy (xapproachesinfinity):

now you can integrate that one

OpenStudy (anonymous):

thanks a lot

OpenStudy (xapproachesinfinity):

welcome!

OpenStudy (anonymous):

sqrt of (1 +sinx)

OpenStudy (xapproachesinfinity):

you mean the result?

OpenStudy (anonymous):

the result after integrating this

OpenStudy (xapproachesinfinity):

doesn't look to me a good one! what did you do?

OpenStudy (xapproachesinfinity):

i haven't done the integral so i don't really know what it would look like

OpenStudy (xapproachesinfinity):

but i bet there should be an ln(...)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} \sqrt{(1 +\sin \theta)}\]

OpenStudy (anonymous):

okiee thanks again

OpenStudy (xapproachesinfinity):

where did you come up with the interval? was it not indefinite integral?

OpenStudy (anonymous):

no...this is the main question

OpenStudy (xapproachesinfinity):

i don't got what you are telling hehe what does that have to do with our integral?

OpenStudy (anonymous):

its a new question...where i have problem...and this one is not a indefinite integral...but i just wanna know that how to do the integration

OpenStudy (xapproachesinfinity):

how about u substitution?

OpenStudy (anonymous):

i tried...but couldn't continue

OpenStudy (xapproachesinfinity):

what u sub did you choose?

OpenStudy (anonymous):

u =sin theta

OpenStudy (xapproachesinfinity):

hmmm... i tried u sub for \(1+sin\theta\) the same problem you have to consider by part

OpenStudy (xapproachesinfinity):

wolfram result http://www.wolframalpha.com/input/?i=%5Cint+sqrt%281%2Bsinx%29dx haha, but i don't think it is that way

OpenStudy (anonymous):

i didn't got a bit :/

OpenStudy (xapproachesinfinity):

haha, you sure that's the real expression?

OpenStudy (anonymous):

yeshh its the real expression...got more weirder integration in stock :/

OpenStudy (xapproachesinfinity):

well, let's stick with u sub haha u=sintheta

OpenStudy (xapproachesinfinity):